# Large Language Model-Brained GUI Agents: A Survey

### v6

@@ 7.1.2 Collection Pipeline
TABLE X: Overview of datasets for optimizing LLMs tailored for GUI agents (Part I).
@@
TABLE XI: Overview of datasets for optimizing LLMs tailored for GUI agents (Part II).
@@
TABLE XII: Overview of datasets for optimizing LLMs tailored for GUI agents (Part III).
@@
7.2 Web Agent Data
Web-based GUI agents demand datasets that capture the intricate complexity and diversity of real-world web interactions. These datasets often encompass varied website structures, including DOM trees and HTML content, as well as multi-step task annotations that reflect realistic user navigation and interaction patterns. Developing agents that can generalize across different websites and perform complex tasks requires comprehensive datasets that provide rich contextual information.
7.3 Mobile Agent Data
Mobile platforms are critical for GUI agents due to the diverse range of apps and unique user interactions they involve. To develop agents that can effectively navigate and interact with mobile interfaces, datasets must offer a mix of single and multi-step tasks, focusing on natural language instructions, UI layouts, and user interactions.
7.4 Computer Agent Data
In contrast to mobile and web platforms, the desktop domain for GUI agents has relatively fewer dedicated datasets, despite its critical importance for applications like productivity tools and enterprise software. However, notable efforts have been made to support the development and evaluation of agents designed for complex, multi-step desktop tasks.
@@ 7.5 Cross-Platform Agent Data
Cross-platform datasets play a pivotal role in developing versatile GUI agents that can operate seamlessly across mobile, computer, and web environments. Such datasets support generalizability and adaptability, enabling agents to handle varied interfaces and tasks in real-world applications.
@@ 8 Models for Optimizing LLM-Brained GUI Agents
TABLE XIII: Overview of foundation models for LLM-brained GUI agents.
@@ 8.1 Foundation Models
Foundation models serve as the core of LLM-powered GUI agents, providing the essential capabilities for understanding and interacting with graphical user interfaces. Recent advancements in both close-source and open-source MLLMs have significantly enhanced the potential of GUI agents, offering improvements in efficiency, scalability, and multimodal reasoning. This subsection explores these foundation models, highlighting their innovations, contributions, and suitability for GUI agent applications. For a quick reference, Table XIII presents an overview of the key models and their characteristics.
@@ 8.2 Large Action Models
In the following sections, we present an analysis of LAMs tailored for GUI agents across different platforms, summarized in Tables XIV, XV, and XVI, followed by an in-depth discussion in the subsequent subsections.
TABLE XIV: An overview of GUI-optimized models (Part I).
@@
TABLE XVI: An overview of GUI-optimized models (Part III).
@@
Omni-Parser [170] Mobile, desktop, and webpage BLIP-2 [293], YOLOv8 [377] - UI screenshots with bounding boxes and numeric IDs, and structured representations including text from OCR and descriptions of functionality IDs, bounding boxes, and descriptions for interactable elements 67,000 UI screenshots labeled with bounding boxes and 7,185 icon-description pairs using GPT-4 Introduces a vision-only screen parsing framework, enabling general UI understanding without reliance on external information, significantly improving action prediction accuracy for LLM-powered agents https://github.com/microsoft/OmniParser
@@
8.3 LAMs for Web GUI Agents
In the domain of web-based GUI agents, researchers have developed specialized LAMs that enhance interaction and navigation within web environments. These models are tailored to understand the complexities of web GUIs, including dynamic content and diverse interaction patterns.
8.4 LAMs for Mobile GUI Agents
Mobile platforms present unique challenges for GUI agents, including diverse screen sizes, touch interactions, and resource constraints. Researchers have developed specialized LAMs to address these challenges, enhancing interaction and navigation within mobile environments.
8.5 LAMs for Computer GUI Agents
For desktop and laptop environments, GUI agents must handle complex applications, multitasking, and varied interaction modalities. Specialized LAMs for computer GUI agents enhance capabilities in these settings, enabling more sophisticated task execution.
@@ 8.6 Cross-Platform Large Action Models
To achieve versatility across various platforms, cross-platform LAMs have been developed, enabling GUI agents to operate seamlessly in multiple environments such as mobile devices, desktops, and web interfaces.
@@ 9.1 Evaluation Metrics
The implementation of metrics in each GUI agent benchmark might vary depending on the platform and the task formulation. In Tables XVII to XXI, we mapped the original metrics used in the benchmarks, which may possess different names, to the categories that we defined above.
@@ 9.3 Evaluation Platforms
All the metrics, measurements, and platforms discussed are essential for a comprehensive evaluation of GUI agents across multiple aspects. Most existing benchmarks rely on them for evaluation. In what follows, we first provide an overview of these benchmarks for GUI agents in Tables XVII to XXI, and then detail them selectively in the following subsections.
TABLE XVII: Overview of GUI agent benchmarks (Part I).
@@
TABLE XVIII: Overview of GUI agent benchmarks (Part II).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
TABLE XIX: Overview of GUI agent benchmarks (Part III).
@@
TABLE $X$: Overview of GUI agent benchmarks (Part IV).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
OSWorld [317] Linux, Windows, macOS, Web 2024 Scalable, real computer environment for multimodal agents, supporting task setup, execution-based evaluation, and interactive learning across Ubuntu, Windows, and macOS. 369 Ubuntu tasks, 43 Windows tasks Task Success Rate Execution-based State Information (such as internal file interpretation, permission management) https://os-world.github.io/
OmniACT [318] MacOS, Linux, Windows, Web 2024 Assesses agents’ capability to generate executable programs for computer tasks across desktop and web applications in various OS environments, prioritizing multimodal challenges. 9,802 data points Task Success Rate, Step Success Rate Action Match https://huggingface.co/datasets/Writer/omniact
TABLE XXI: Overview of GUI agent benchmarks (Part V).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
VideoGUI [319] Windows, Web 2024 Focuses on visual-centric tasks from instructional videos, emphasizing planning and action precision in applications like Adobe Photoshop and Premiere Pro. 178 tasks, 463 subtasks Task Success Rate State Information, Action Match https://showlab.github.io/videogui
9.4 Web Agent Benchmarks
Evaluating GUI agents in web environments necessitates benchmarks that capture the complexities and nuances of web-based tasks. Over the years, several benchmarks have been developed, each contributing unique perspectives and challenges to advance the field.
@@ 9.5 Mobile Agent Benchmarks
Evaluating GUI agents on mobile platforms presents unique challenges due to the diversity of interactions and the complexity of mobile applications. Several benchmarks have been developed to address these challenges, each contributing to the advancement of mobile agent evaluation.
@@ 9.6 Computer Agent Benchmarks
Evaluating GUI agents on desktop computers involves diverse applications and complex workflows. Several benchmarks have been developed to assess agents’ capabilities in these environments, each addressing specific challenges and advancing the field.
@@ 9.7 Cross-Platform Agent Benchmarks
To develop GUI agents capable of operating across multiple platforms, cross-platform benchmarks are essential. These benchmarks challenge agents to adapt to different environments and interfaces, evaluating their versatility and robustness.
@@ 10.1 GUI Testing
TABLE XXII: Overview of GUI-testing with LLM-powered GUI agents.
@@
Figure 25 illustrates the use of an LLM-powered GUI agent to test font size adjustment on Windows OS. With only a natural language test case description, the agent autonomously performs the testing by executing UI operations, navigating through the settings menu, and leveraging its screen understanding capabilities to verify the final outcome of font size adjustment. This approach dramatically reduces the effort required for human or script-based testing. Next, we detail the GUI testing works powered by GUI agents, and first provide an overview Table XXII.
@@ 10.2 Virtual Assistants
TABLE XXIII: Overview of virtual assistants with LLM-powered GUI agents.
@@
To explore more real-world applications of virtual assistants powered by GUI agents, we provide an overview of advancements across research, open-source initiatives, and production-level applications, as summarized in Table XXIII.

### v7

@@ 7.1.2 Collection Pipeline
TABLE XIII: Overview of datasets for optimizing LLMs tailored for web GUI agents.
@@
7.2 Web Agent Data
Web-based GUI agents demand datasets that capture the intricate complexity and diversity of real-world web interactions. These datasets often encompass varied website structures, including DOM trees and HTML content, as well as multi-step task annotations that reflect realistic user navigation and interaction patterns. Developing agents that can generalize across different websites and perform complex tasks requires comprehensive datasets that provide rich contextual information. We provide an overview of web-based GUI agents dataset in Table XIII.
TABLE XIV: Overview of datasets for optimizing LLMs tailored for mobile GUI agents (Part I).
@@ 7.2 Web Agent Data
TABLE XV: Overview of datasets for optimizing LLMs tailored for mobile GUI agents (Part II).
@@
7.3 Mobile Agent Data
Mobile platforms are critical for GUI agents due to the diverse range of apps and unique user interactions they involve. To develop agents that can effectively navigate and interact with mobile interfaces, datasets must offer a mix of single and multi-step tasks, focusing on natural language instructions, UI layouts, and user interactions. We first overview mobile GUI agents dataset in Tables XIV and XV.
TABLE XVI: Overview of datasets for optimizing LLMs tailored for computer GUI agents.
@@ 7.3 Mobile Agent Data
7.4 Computer Agent Data
In contrast to mobile and web platforms, the desktop domain for GUI agents has relatively fewer dedicated datasets, despite its critical importance for applications like productivity tools and enterprise software. However, notable efforts have been made to support the development and evaluation of agents designed for complex, multi-step desktop tasks. We show related dataset for computer GUI agents in Table XVI.
TABLE XVII: Overview of datasets for optimizing LLMs tailored for cross-platform GUI agents.
@@ 7.4 Computer Agent Data
OS-Genesis [285] Web and Android Reverse task synthesis, where the GUI environment is explored interactively without predefined tasks or human annotations. High-level instructions, low-level instructions, action sequences, and environment states. 1,000 synthesized trajectories. Model-based interaction-driven approach with GPT-4o. Reverses the conventional task-driven data collection process by enabling exploration-first trajectory synthesis. https://qiushisun.github.io/OS-Genesis-Home/
@@ 7.5 Cross-Platform Agent Data
Cross-platform datasets play a pivotal role in developing versatile GUI agents that can operate seamlessly across mobile, computer, and web environments. Such datasets support generalizability and adaptability, enabling agents to handle varied interfaces and tasks in real-world applications. We provide an overview of related dataset for cross-platform GUI agents in Table XVII.
@@
OS-Genesis [285] adopts a reverse task synthesis approach for the Android and web platforms. It leverages GPT-4o to interactively explore the environment and generate instructions in a reverse manner. This process constructs high-quality, diverse GUI trajectories without relying on human annotations or predefined tasks. By eliminating these dependencies, OS-Genesis achieves scalable and efficient training for GUI agents while significantly enhancing the diversity and quality of the generated data.
@@ 8 Models for Optimizing LLM-Brained GUI Agents
TABLE XVIII: Overview of foundation models for LLM-brained GUI agents.
@@ 8.1 Foundation Models
Foundation models serve as the core of LLM-powered GUI agents, providing the essential capabilities for understanding and interacting with graphical user interfaces. Recent advancements in both close-source and open-source MLLMs have significantly enhanced the potential of GUI agents, offering improvements in efficiency, scalability, and multimodal reasoning. This subsection explores these foundation models, highlighting their innovations, contributions, and suitability for GUI agent applications. For a quick reference, Table XVIII presents an overview of the key models and their characteristics.
@@ 8.2 Large Action Models
TABLE XIX: An overview of GUI-optimized models on web platforms.
@@
8.3 LAMs for Web GUI Agents
In the domain of web-based GUI agents, researchers have developed specialized LAMs that enhance interaction and navigation within web environments. These models are tailored to understand the complexities of web GUIs, including dynamic content and diverse interaction patterns. We present an analysis of LAMs tailored for web GUI agents in Table XIX.
TABLE $X$: An overview of GUI-optimized models on mobile platforms.
@@ 8.3 LAMs for Web GUI Agents
8.4 LAMs for Mobile GUI Agents
Mobile platforms present unique challenges for GUI agents, including diverse screen sizes, touch interactions, and resource constraints. Researchers have developed specialized LAMs to address these challenges, enhancing interaction and navigation within mobile environments. We present an overview of LAMs tailored for mobile GUI agents in Table XX.
TABLE XXI: An overview of GUI-optimized models on computer platforms.
@@ 8.4 LAMs for Mobile GUI Agents
8.5 LAMs for Computer GUI Agents
For desktop and laptop environments, GUI agents must handle complex applications, multitasking, and varied interaction modalities. Specialized LAMs for computer GUI agents enhance capabilities in these settings, enabling more sophisticated task execution. We overview of LAMs for computer GUI agents across in Table XXI.
TABLE XXII: An overview of GUI-optimized models on cross-platform agents.
@@ 8.6 Cross-Platform Large Action Models
To achieve versatility across various platforms, cross-platform LAMs have been developed, enabling GUI agents to operate seamlessly in multiple environments such as mobile devices, desktops, and web interfaces. We provide an analysis of LAMs tailored for cross-platform GUI agents in Table XXII.
@@ 9.1 Evaluation Metrics
The implementation of metrics in each GUI agent benchmark might vary depending on the platform and the task formulation. In all tables in this section, we mapped the original metrics used in the benchmarks, which may possess different names, to the categories that we defined above.
@@ 9.3 Evaluation Platforms
All the metrics, measurements, and platforms discussed are essential for a comprehensive evaluation of GUI agents across multiple aspects. Most existing benchmarks rely on them for evaluation. In what follows, detail these benchmarks for GUI agents selectively.
TABLE XXIII: Overview of web GUI agent benchmarks (Part I).
@@
TABLE XXIV: Overview of web GUI agent benchmarks (Part II).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
TABLE XXV: Overview of web GUI agent benchmarks (Part III).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
BrowserGym [344] Web 2024 Provides a unified, extensible, and open-source environment for evaluating web agents with standardized APIs and observations. Benchmarks include MiniWoB(++) with 125 tasks, WebArena with 812 tasks, and WorkArena with up to 341 tasks per level. Task Success Rate, Step Success Rate, Turn Success Rate, Efficiency Metrics. Text-based matching and element match. https://github.com/ServiceNow/BrowserGym
WebWalkerQA [345] Web 2025 Benchmarks the capacity of LLMs to handle deep, structured, and realistic web-based navigation and reasoning tasks. 680 high-quality QA pairs. Task Success Rate, Efficiency Score. Text Match, Action Match. https://github.com/Alibaba-NLP/WebWalker
9.4 Web Agent Benchmarks
Evaluating GUI agents in web environments necessitates benchmarks that capture the complexities and nuances of web-based tasks. Over the years, several benchmarks have been developed, each contributing unique perspectives and challenges to advance the field. We first provide an overview of these benchmarks in Tables XXIII, XXIV and XXV.
BrowserGym [344] builds ecosystem designed for web agent research. It unifies various benchmarks like MiniWoB(++) [138], WebArena [327], and WorkArena [337] under a single framework, addressing the issue of fragmentation in web agent evaluation. By leveraging standardized observation and action spaces, it enables consistent and reproducible experiments. BrowserGym’s extensible architecture make it a vital tool for developing and testing GUI-driven agents powered by LLMs, significantly accelerating innovation in web automation research.
TABLE XXVI: Overview of mobile GUI agent benchmarks (Part I).
@@ 9.4 Web Agent Benchmarks
TABLE XXVII: Overview of mobile GUI agent benchmarks (Part II).
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@
GTArena [356] Mobile applications 2024 Introduces a Transition Tuple for GUI defects, enabling large-scale defect dataset creation and reproducible, end-to-end automated testing. 10,000+ GUI display and GUI interactions Task Success Rate, Step Success Rate. Text Match, Element Match, Action Match. https://github.com/ZJU-ACES-ISE/ChatUITest
A3 [357] Mobile Android 2025 Introduces a novel business-level LLM-based evaluation process, significantly reducing human labor and coding expertise requirements. 201 tasks across 21 widely used apps. Task Success Rate. Element Match, Action Match. https://yuxiangchai.github.io/Android-Agent-Arena/
@@
SPHINX [358] Android 2025 Provides a fully automated benchmarking suite and introduces a multi-dimensional evaluation framework. 284 common tasks. Task Success Rate, Efficiency Score, Completion under Policy, Turn Success Rate. Text Match, Image Match, Element Match, Action Match. /
@@ 9.5 Mobile Agent Benchmarks
Evaluating GUI agents on mobile platforms presents unique challenges due to the diversity of interactions and the complexity of mobile applications. Several benchmarks have been developed to address these challenges, each contributing to the advancement of mobile agent evaluation. We first provide an analysis for these mobile benchmarks in Tables XXVI and XXVII.
@@
Zhao et al., introduce GTArena [356], a formalized framework and benchmark designed to advance autonomous GUI testing agents. GTArena provides a standardized evaluation environment tailored for multimodal large language models. Central to its design is the novel Transition Tuple data structure, which systematically captures and analyzes GUI defects. The benchmark assesses three core tasks—test intention generation, task execution, and defect detection—using a diverse dataset comprising real-world, artificially injected, and synthetic defects, establishing GTArena as a pioneering benchmark for GUI testing agents.
@@
TABLE XXVIII: Overview of computer GUI agent benchmarks.
Benchmark Platform Year Highlight Data Size Metric Measurement Link
OSWorld [334] Linux, Windows, macOS 2024 Scalable, real computer environment for multimodal agents, supporting task setup, execution-based evaluation, and interactive learning across Ubuntu, Windows, and macOS. 369 Ubuntu tasks, 43 Windows tasks Task Success Rate Execution-based State Information (such as internal file interpretation, permission management) https://os-world.github.io/
OmniACT [335] MacOS, Linux, Windows 2024 Assesses agents’ capability to generate executable programs for computer tasks across desktop and web applications in various OS environments, prioritizing multimodal challenges. 9,802 data points Task Success Rate, Step Success Rate Action Match https://huggingface.co/datasets/Writer/omniact
VideoGUI [336] Windows 2024 Focuses on visual-centric tasks from instructional videos, emphasizing planning and action precision in applications like Adobe Photoshop and Premiere Pro. 178 tasks, 463 subtasks Task Success Rate State Information, Action Match https://showlab.github.io/videogui
@@ 9.6 Computer Agent Benchmarks
Evaluating GUI agents on desktop computers involves diverse applications and complex workflows. Several benchmarks have been developed to assess agents’ capabilities in these environments, each addressing specific challenges and advancing the field. We overview benchmarks for computer GUI agents in Table XXVIII.
@@
TABLE XXIX: Overview of cross-platform GUI agent benchmarks.
Benchmark Platform Year Highlight Data Size Metric Measurement Link
@@ 9.7 Cross-Platform Agent Benchmarks
To develop GUI agents capable of operating across multiple platforms, cross-platform benchmarks are essential. These benchmarks challenge agents to adapt to different environments and interfaces, evaluating their versatility and robustness. We provide an overview of benchmarks for cross-platform GUI agents in Tables XXIX.
@@ 10.1 GUI Testing
TABLE XXX: Overview of GUI-testing with LLM-powered GUI agents (Part I).
@@
TABLE XXXI: Overview of GUI-testing with LLM-powered GUI agents (Part II).
@@
Figure 25 and illustrates the use of an LLM-powered GUI agent to test font size adjustment on Windows OS. With only a natural language test case description, the agent autonomously performs the testing by executing UI operations, navigating through the settings menu, and leveraging its screen understanding capabilities to verify the final outcome of font size adjustment. This approach dramatically reduces the effort required for human or script-based testing. Next, we detail the GUI testing works powered by GUI agents, and first provide an overview Tables XXX and XXXI.
@@ 10.2 Virtual Assistants
TABLE XXXII: Overview of virtual assistants with LLM-powered GUI agents (Part I).
@@
TABLE XXXIII: Overview of virtual assistants with LLM-powered GUI agents (Part II).
@@
Eko [424] Production Web browsers and computer environments ChatGPT and Claude 3.5 Visual-Interactive Element Perception (VIEP) technology for interacting with GUI elements. Standard UI operations. Automates tasks by handling diverse workflows. Decomposes natural language task descriptions into executable workflows, enabling seamless integration of natural language and programming logic in agent design. https://eko.fellou.ai/
@@
To explore more real-world applications of virtual assistants powered by GUI agents, we provide an overview of advancements across research, open-source initiatives, and production-level applications, as summarized in Table XXXII and XXXIII.
@@ 10.2.3 Production
Eko [424] serves as a prime example of a versatile and efficient tool for developing intelligent agents capable of interacting with GUIs across various platforms. Its integration with multiple LLMs and the innovative Visual-Interactive Element Perception (VIEP) technology highlight its capability to perform complex tasks through natural language instructions. Eko’s comprehensive tool support make it a valuable resource for developers aiming to create customizable and production-ready agent-based workflows. By facilitating seamless interaction within GUI environments, Eko exemplifies the advancements in virtual assistants powered by LLMs.